OpenStudy (anonymous):

Find the solution to the system of equations. x+y+z=11 y+z=12 z=8

OpenStudy (anonymous):

y + 8 = 12 --> y = 4 x + 4 + 8 = 11 --> x = -1

OpenStudy (gw2011):

y+z=12 y+8=12 y=4 x+y+z=11 x+4+8=11 x+12=11 x=-1